OpenStudy (anonymous):

A 12 ft. ladder leaning against a wall has its base 4 ft. from the wall. Find the sine and the cosine of the angle that the ladder makes with the ground.

OpenStudy (anonymous):



OpenStudy (anonymous):

you need the other side of the triangle, which you find via pythagoras

OpenStudy (anonymous):

it is \(\sqrt{12^2-4^2}\) once you find it, find the sine and cosine using the appropriate ratios

OpenStudy (anonymous):

I just typed that formula and the answer I got is -4. :(

OpenStudy (anonymous):

some mistake there

OpenStudy (anonymous):

\(\sqrt{1^2-4^2}=\sqrt{144-16}=\sqrt{128}\)

OpenStudy (radar):

Now that you have all three side you can solve. Note that you can solve for the cosine with the facts given: Cos x= 4/12= 1/3 You can now solve for the sine using the identity :